Carla Sagers sells three capital assets this year. The first is stock she received as a bequest from her grandmother early in 2009. Her grandmother’s basis was $2,000, but the stock’s FMV upon her death was $30,000, and Carla sold it for $28,000. The second asset Carla sold was rental property she held as community property with her late husband until he died in 2011. They bought it in 1992 for $85,000, and it was appraised in 2011 for $105,000. Carla was able to sell it for $107,000, after selling expenses. The third asset was a gift of rare coins from her uncle. He bought them in 1978 for $20,000 and gave them to Carla in 2000, when they were worth $30,000. Her uncle paid gift tax on the transfer in the amount of $4,500 (her uncle had given her a gift equal to the gift tax annual exclusion earlier that year). Carla sold the coins for $31,000. What is Carla’s net gain on these three transactions?

Question 19 options:

$19,500.

$9,500.

$6,500.

$1,000.
Answer #1

Net loss on selling Grandma's stock = 30000-28000 = -2000

The asset value was $30000 but she sold it for less.

Net profit on rental property =$107000-105000 = $2000

We consider the appraised value of property.

Now , Net gain on selling coins = Difference in value at time of selling= 31000-30000 = $1000

Total net gain = $1000